hey i have another problem for you to help me out with if you could?
F street has an equation of y=x+4. Determine the x-intercept and y-intercept of f street.
thanks ALOT!

Answer by 303795(602)

The x intercept is when y = 0 and the y intercept is when x = 0.
Substitute y=0 into the equation to find the x intercept.
y=x+4
0=x+4
x=-4 (x intercept)
Substitute x=0 into the equation to find the y intercept.
y=x+4
y=0+4
y=4 (y intercept)